For the following three vectors, what is 2 . c . ?A × B | ? A 3.00i 3.00j 3.00k B 4.00i+ 3.00j+ 3.00k C = 7.00;-9.00/ Number Units
Answer #1

2A×B)=16.00 6.00-6.00 (2.A x B) (6x3)-(-6x3))-j(6x3)-(-4x-6)+(6x3)-(-4x6)) 2A× B) = i(18 + 18)-3(18-24) + k(18 + 24) 一4.00 3.